n the market for cashews the following events are taking place. Explain the effect of each single event on demand, supply, the equilibrium price and the equilibrium quantity for parts a)-c), and then answer the rest of the questions.

a) The price of hazelnuts, a substitute in consumption, increases. Explain which curve will shift, and how the equilibrium price and quantity will change. 

b) A hailstorm destroys 50% of all cashew trees. 

c) For Sam, cashews are an inferior good. Suppose that Sam's income increases, because he gets a promotion. At the same time, there are more cashew plantations starting up.

d) When the price of a pound of cashews increases from $30 to $35, consumption drops from 1000 to 500 pounds. Use the midpoint formula to calculate the price elasticity of demand coefficient. Is demand elastic or inelastic?

e) Given your results from part d) would revenues for farmers increase or decrease with this increase in prices? How does this relate to elasticity? Are consumers very responsive to this price change?

f) Why would the demand for cashews be inelastic for some individuals? Explain, incorporating one of the determinants of price elasticity in your explanation.


1
Expert's answer
2021-03-04T15:09:51-0500

(a) Demand curve will shift to the right because the demand for cashew will increase. The equilibrium price will increase as the equilibrium quantity decreases.

(b) Demand remaining the same, the supply curve will shift to the left ,equilibrium price will rise as the equilibrium quantity decreases.

(c) The supply curve will shift to the right as the demand curve shifts to the left. Both equilibrium price and quantity will decrease.

(d)"e_{(p)}=\frac{dQ/Q}{dP/P}"

"=(1000+500)\/2=750"

"=(500-1000)\/750=-0.67"

"=67" %

"=(30+35)\/2=32.5"

"=(35-30)\/32.5=0.15"

"=15%"

"=67\/15=4.47"

"=-4.47"

The demand is inelastic.

(e)The revenue will increase. Inelastic goods increase prices thereby selling a little less but making higher revenues. Consumers are responsible for the change in price due to their change in demand for the good.

(f) The demand for cashews will be inelastic for some individuals because the goods which are price inelastic tend to have fewer substitutes and are considered to be necessities by users.